OpenStudy (anonymous):

Find an equation of a line through (10, −10) and parallel to 20x − y = 2.

OpenStudy (anonymous):

If a line is parallel to another line, it has the same slope as the other line.

OpenStudy (anonymous):

The slope of 20x − y = 2 is -A/B = -20/-1 = 20. Then plug that and the one point into the point-slope form to get: (y + 10) = 20(x - 10) y = 20x - 200 - 10 y = 20x - 210

OpenStudy (anonymous):

Thanks so much @ASNinja

OpenStudy (anonymous):

np